OpenStudy (anonymous):

Simplify 12 times x to the fifth power all over 20 times x squared.

OpenStudy (jackellyn):

12x^5/20x^2 is this the problem?

OpenStudy (anonymous):

yes

OpenStudy (jackellyn):

Well first divide the coefficients. 12/20 = what?

OpenStudy (anonymous):

1.6 ?

OpenStudy (jackellyn):

12 2^2 x 3 -- -------- 20 2^2 x 5 You can can cancel the 2^2 which leaves you with 3 - 5 so 12 3 -- = -- 20 5

OpenStudy (jackellyn):

Next step, when you're dividing variables, the exponents subtract. So, x^5 --- = 5-2 = what? x^2

OpenStudy (anonymous):

3

OpenStudy (jackellyn):

so x^5 --- = x^3 x^2 Therefore, 12x^5 3x^3 ------ = ---- 20x^2 5

OpenStudy (anonymous):

thank you. i hate math so much.. youre a major help.